Line L has equation y= -5x-15. Line m has equation y=x-3. Find the point of intersection. Can you also please show your work to get your answer.
Answer by mananth(16949)

y= -5x-15.
y=x-3
Find the Y & X intercepts of the two lines by plugging x=0, & y=0
equation1
when x= 0, y= -15 ( 0, , -15 )
when y = 0, x= -3 ( -3 , 0, )
Equation 2
when x= 0, y= -3 ( 0, , -3 )
when y = 0, x= 3 ( 3 , 0, )

the solution is (-2,-5)